OpenStudy (anonymous):

A church steeple casts a shadow 114ft long, and same time a 8.0ft post casts a shadow 5.0ft long. How high is the steeple? Round to the nearest cent.

OpenStudy (anonymous):

182.4

OpenStudy (anonymous):

how'd u do that?

OpenStudy (anonymous):

im not 100% sure it is correct but i used a ratio of 8/5=x/114 and solved for x

OpenStudy (anonymous):

it's wrong....

OpenStudy (anonymous):

Solve the following for x:\[\frac{8}{5}=\frac{x}{114}\]\[x=\frac{912}{5}=182.4 \]